For each of the following strong base solutions, determine [OH−],[H3O+] , pH, and pOH.

0.19 MM NaOH

1.3×10−3 MM Ca(OH)2

4.7×10−4 MM Sr(OH)2

8.6×10−5 MM KOHKOH
Answer #1

- nat toll 0 0.19 m Nuolt Naoh Coll] = 0-13 m kwe (1438t) = x [oli) = 1,514 47 (M430) = [1430+) = level 2([142043 = 5-26410